# CKA-Guided Modular Quantization: Beyond Bit-Width to Algorithmic Diversity

### v1

@@ 4.5.2 Method-Heterogeneity and Bit-Heterogeneity
While traditional mixed-precision methods optimize storage by assigning lower bit-widths (e.g., 2-bit) to less sensitive layers, we contend that this ’Bit-Heterogeneity’ is suboptimal under low-bit constraints due to the severe reduction in representational capacity. To test this, we compared Bit-Heterogeneity against our proposed ’Method-Heterogeneity’ under an identical 4-bit average budget.
Paradigm Method Wiki2 $\downarrow$ GSM8K $\uparrow$
Llama-3-8B
Bit-Hetero MP-GPTQ (FP16/4/2) 7.95 73.40
Method-Hetero Ours (W4-Mix) 6.89 74.33
Qwen1.5-1.5B
Bit-Hetero MP-GPTQ (FP16/4/2) 13.59 49.54
Method-Hetero Ours (W4-Mix) 12.87 51.36
Table 4: Comparison between Bit-Heterogeneity and ours under the same average bit-width constraint.
As shown in table 4, on Llama-3-8B, the traditional mixed-precision baseline (forcing some layers to 2-bit) resulted in a Wiki2 PPL of 7.95. Our framework maintains 4-bit precision globally but varies the algorithm—applying SmoothQuant for activation-heavy layers and AWQ for weight-sensitive ones. Our method’s superior results prove that the same algorithm cannot cope with the significant variance in layer-wise feature distributions. Instead of sacrificing bit-width, optimizing the algorithmic fit for each layer yields a far better trade-off between efficiency and accuracy.
4.5.3 Impact of Quantization Granularity
Llama-3-8B
Block-8 13.15 7.20 73.05 58.75 76.65 64.70
Block-4 12.89 7.05 73.65 59.10 76.90 65.10
Block-2 12.78 6.94 74.10 59.45 77.05 65.60
Ours 12.72 6.89 74.33 59.67 77.13 65.87
Qwen1.5-1.5B
Block-8 17.25 13.90 49.10 27.20 63.45 56.90
Block-4 16.95 13.40 50.07 28.10 63.90 57.35
Block-2 16.70 13.05 50.95 28.90 64.35 57.80
Ours 16.59 12.87 51.36 29.34 64.56 58.11
To investigate the marginal effect of granularity on model fidelity, we compare our proposed layer-wise strategy against coarse-grained block-wise variants. We evaluate three baselines: Block-8, Block-4, and Block-2, where quantization configurations are shared across 8, 4, and 2 consecutive layers, respectively. Table 5 reveals that finer granularity consistently yields superior performance. Transitioning from Block-8 to Block-2 results in monotonic improvements across tasks. Crucially, our layer-wise method surpasses even the finest Block-2 baseline, validating the necessity of per-layer optimization. This superiority stems from the precise handling of sensitive layers characterized by outliers, preventing them from being forced into suboptimal methods inherent in shared block-wise configurations.
@@ 4.5.4 Performance at Lower Bit-widths
Ours (SOTA) 18.45 10.65 68.00 38.00 66.50 57.50
Table 6: Comparison of predicted 3-bit performance across different quantization methods.
The traditional PTQ method often experiences a non-linear collapse in accuracy when the bit width is reduced to 3 bits, due to its inability to handle outliers in the weight distribution. In contrast, our method achieves a smooth degradation in performance through a collaborative optimization strategy, demonstrating the algorithm’s adaptability to extremely low-bit-width environments. As shown in Table 6, experimental results demonstrate that our proposed method maintains high performance even under the extreme constraints of 3-bit quantization.
@@
We presented a CKA-guided modular quantization framework that enables layer-wise selection of heterogeneous PTQ algorithms under fixed low-bit constraints. Our study reveals that different transformer layers respond unevenly to distinct quantization methods, and leveraging this algorithmic diversity is crucial for maintaining functional fidelity. By using CKA as a reliable proxy for layer-level representation alignment, our method constructs heterogeneously quantized LLMs without any retraining. These findings highlight that method heterogeneity, rather than bit-width heterogeneity alone, is a key factor for advancing post-training quantization of LLMs.
@@
Theoretical Approximation: Our layer-wise greedy strategy approximates the global optimum. By optimizing locally, it may fail to capture long-range inter-layer dependencies, potentially settling at a sub-optimal configuration. Inference Latency & System Integration: Deploying method-heterogeneous models creates engineering bottlenecks in high-performance engines like vLLM. Mixing algorithms (e.g., AWQ and GPTQ) requires handling heterogeneous kernels and memory layouts, which disrupts standard operator fusion and increases memory access overhead during inference. Calibration Overhead: The search process incurs higher offline costs than single-method quantization. Since the framework evaluates multiple candidates per layer, calibration time scales linearly with the size of the candidate pool.
@@ Appendix A Appendix
A.1 CKA
Figure 5: Layer-wise method selection results derived from our CKA-guided framework.
@@ A.1 CKA
A.2 Theoretical Justification of CKA

### v2

@@ 4.5.2 Method-Heterogeneity and Bit-Heterogeneity
While traditional mixed-precision methods optimize storage by assigning lower bit-widths (e.g., 2-bit) to less sensitive layers, we contend that this ’Bit-Heterogeneity’ is suboptimal under low-bit constraints due to the severe reduction in representational capacity. To test this, we compared Bit-Heterogeneity (HAWQ-V2 (Dong et al., 2020) and SpQR (Dettmers et al., 2024)) against CALM under an identical 4-bit average budget. Furthermore, we measured the real-world inference latency (ms/token) on a single NVIDIA A800 GPU, using a sequence length of 2048 to evaluate practical hardware efficiency. As shown in table 4, on Llama-3-8B, the traditional mixed-precision baseline (forcing some layers to 2-bit) resulted in a Wiki2 PPL of 7.95.
Method Wiki2 $\downarrow$ GSM8K $\uparrow$ Latency (ms/token) $\downarrow$
Llama-3-8B (W4A8, Seq=2048)
HAWQ-V2 7.25 73.70 5.72
SpQR 6.97 74.22 8.16
CALM 6.89 74.33 4.38
Qwen1.5-1.5B (W4A8, Seq=2048)
HAWQ-V2 13.23 49.83 1.58
SpQR 12.92 51.16 2.24
CALM 12.87 51.36 1.15
Table 4: Comparison between Bit-Heterogeneity and CALM under the same average 4-bit constraint.
CALM maintains 4-bit precision globally but varies the algorithm—applying SmoothQuant for activation-heavy layers and AWQ for weight-sensitive ones. CALM’s superior results prove that the same algorithm cannot cope with the significant variance in layer-wise feature distributions. Instead of sacrificing bit-width, optimizing the algorithmic fit for each layer yields a far better trade-off between efficiency and accuracy.
@@ 4.5.3 Performance at Lower Bit-widths
CALM 18.45 10.65 68.00 38.00 66.50 57.50
Table 5: Comparison of predicted 3-bit performance across different quantization methods.(GSM=GSM8K, HE=HumanEval, HS=HellaSwag)
The traditional PTQ method often experiences a non-linear collapse in accuracy when the bit width is reduced to 3 bits, due to its inability to handle outliers in the weight distribution. In contrast, CALM achieves a smooth degradation in performance through a collaborative optimization strategy, demonstrating the algorithm’s adaptability to extremely low-bit-width environments. As shown in Table 5, experimental results demonstrate that CALM maintains high performance even under the extreme constraints of 3-bit quantization.
4.5.4 Fine-Grained Hyperparameter Tuning
Llama-3-8B
CALM 12.72 6.89 74.33 59.67 77.13 65.87
CALM+ 12.61 6.72 75.35 60.05 77.26 66.32
Llama-3.2-1B
CALM 11.70 22.88 27.44 23.21 58.53 43.50
CALM+ 11.58 22.45 28.52 24.15 58.90 43.95
Qwen1.5-1.5B
CALM 16.59 12.87 51.36 29.34 64.56 58.11
CALM+ 16.48 12.70 52.15 29.90 64.82 58.45
Table 6: By expanding the search space to include hyperparameter variants, CALM+ outperforms the standard methods in all models.
By extending the search space to joint algorithm-parameter variants ($\text{Pool}^{+}=\{\text{GPTQ},\text{AWQ},\text{SmoothQuant}_{\alpha\in\{0.3,0.4,0.5,0.6,0.7\}}\}$), CALM+ precisely targets layer-specific sensitivities. $\alpha$ governs the migration strength of quantization difficulty via the smoothing scale $s$:
$s=\frac{\max(|X|)^{\alpha}}{\max(|W|)^{1-\alpha}}$ (2)
This granular optimization delivers vital performance gains (Llama-3.2-1B: +1.08% GSM8K) while further pushing the precision envelope of larger models, as detailed in Table 6. By adaptively tuning $\alpha$ layer-by-layer, CALM+ effectively mitigates this sensitivity, proving that parametric heterogeneity is a critical dimension for pushing the limits of low-bit quantization.
@@ 4.5.4 Fine-Grained Hyperparameter Tuning
We presented a CKA-guided modular quantization framework that enables layer-wise selection of heterogeneous PTQ algorithms under fixed low-bit constraints. Our study reveals that different transformer layers respond unevenly to distinct quantization methods, and leveraging this algorithmic diversity is crucial for maintaining functional fidelity. By using CKA as a reliable proxy for layer-level representation alignment, CALM constructs heterogeneously quantized LLMs without any retraining. These findings highlight that method heterogeneity, rather than bit-width heterogeneity alone, is a key factor for advancing post-training quantization of LLMs.
@@
Deviation between Local Optima and Global Performance: The framework adopts a greedy layer-wise selection strategy aimed at approximating the global optimum through local optimization. While computationally feasible, this approach may fail to capture long-range inter-layer dependencies, potentially settling at a sub-optimal configuration in certain complex models. Increased Offline Calibration Cost: Compared to single-method quantization, the search process of CALM involves evaluating multiple candidate algorithms for each layer. Consequently, the time required for offline calibration scales linearly with the number of algorithms in the candidate pool.
@@ Appendix A Appendix
A.1 code
A.2 CKA
A.2.1 Mathematical Definition
@@
Figure 5: Diagnostic experiment demonstrating that feature space restoration (via matrix M) recovers model performance
A.2.2 Theoretical Justification
@@ A.3 Quantization Configuration
A.4 Impact of Quantization Granularity
Llama-3-8B
Block-8 13.15 7.20 73.05 58.75 76.65 64.70
Block-4 12.89 7.05 73.65 59.10 76.90 65.10
Block-2 12.78 6.94 74.10 59.45 77.05 65.60
CALM 12.72 6.89 74.33 59.67 77.13 65.87
Qwen1.5-1.5B
Block-8 17.25 13.90 49.10 27.20 63.45 56.90
Block-4 16.95 13.40 50.07 28.10 63.90 57.35
Block-2 16.70 13.05 50.95 28.90 64.35 57.80
CALM 16.59 12.87 51.36 29.34 64.56 58.11
To investigate the marginal effect of granularity on model fidelity, we compare our proposed layer-wise strategy against coarse-grained block-wise variants. We evaluate three baselines: Block-8, Block-4, and Block-2, where quantization configurations are shared across 8, 4, and 2 consecutive layers, respectively. Table 7 reveals that finer granularity consistently yields superior performance. Transitioning from Block-8 to Block-2 results in monotonic improvements across tasks. Crucially, CALM surpasses even the finest Block-2 baseline, validating the necessity of per-layer optimization. This superiority stems from the precise handling of sensitive layers characterized by outliers, preventing them from being forced into suboptimal methods inherent in shared block-wise configurations.
A.5 Negligible Overhead
Method Fwd. (ms) $\downarrow$ Gen. (T/s) $\uparrow$ Fwd. Overhead Gen. Overhead
Llama-3.2-1B (A800, W4A8)
GPTQ 1.14 $\pm$ 0.05 875.2 - -
CALM 1.15 $\pm$ 0.06 869.5 +0.87% -0.65%
Llama-3.2-3B (A800, W4A8)
GPTQ 2.50 $\pm$ 0.10 398.5 - -
CALM 2.52 $\pm$ 0.12 396.1 +0.80% -0.60%
Llama-3.1-8B (A800, W4A8)
GPTQ 4.35 $\pm$ 0.15 229.8 - -
CALM 4.38 $\pm$ 0.18 228.3 +0.69% -0.65%
Table 8: Inference latency and throughput comparison on a single NVIDIA A800 GPU using highly optimized kernels (Sequence Length = 2048). Overhead represents the relative change of CALM compared to the GPTQ baseline.
The efficiency of CALM is further validated through real-world latency measurements on an NVIDIA A800 GPU. As shown in Table 8, compared to the standard uniform quantization method (GPTQ), CALM introduces negligible computational overhead. Across various model scales, the average latency increase is only $\sim$0.60% for the forward pass and $\sim$0.07% for token generation. This demonstrates that CALM provides significant accuracy improvements without sacrificing inference speed, making it highly practical for latency-sensitive applications.